OpenStudy (anonymous):

Put the equation below in the form ax2 + bx + c = 0. Enter exponents using the caret ( ^ ). For example, you would enter 4x2 as 4x^2. -x2 + 4 = (3x - 1)2

OpenStudy (ajprincess):

-x^2+4=(3x-1)^2 Is this the question?

OpenStudy (anonymous):

yes.

OpenStudy (ajprincess):

Can u tel me vat do u get ven u expand (3x-1)^2 plz?

OpenStudy (anonymous):

i dont know how to do any of it.

OpenStudy (anonymous):

im taking a quiz right now am im stuck on this.

OpenStudy (ajprincess):

Ok no problem. I will guide u. (a-b)^2=a^2-2ab+b^2 Here a=3x and b=1 Nw can u find?

OpenStudy (anonymous):

i got 9x^2-1=3x^2-6x+1?

OpenStudy (ajprincess):

It is 9x^2-6x+1. Nw -x^2+4= 9x^2-6x+1. 9x^2+x^2-6x+1-4=0 Simplify this to get ur answer.

OpenStudy (anonymous):

9x^2+x^2-6x+1-4=0 is the answer?

OpenStudy (ajprincess):

no u need to simplify further.

OpenStudy (anonymous):

10x^2-6x-3=0?

OpenStudy (ajprincess):

yes.

OpenStudy (anonymous):

yay it was right. do you mind helping me with the rest?

OpenStudy (ajprincess):

sure post them.

OpenStudy (ajprincess):

it wud b really better f u post them as new questions.

OpenStudy (anonymous):

x2 - 64 = 0

OpenStudy (anonymous):

it would be 8 and -8 right?

OpenStudy (ajprincess):

yes

OpenStudy (anonymous):

i dont know how to post a new question

OpenStudy (ajprincess):

To post a new question u need to close this question. Below ur questions u will find two buttons. One is bump and the other is close. Click close to close this question. Then u will b able to post a new question.

OpenStudy (anonymous):

k